OpenStudy (anonymous):

Find the slope of the line that passes through (3,5) and (-1,4).

OpenStudy (carson889):

Slope = run over rise. Slope = \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] Thus slope = \[\frac{ 4-5 }{ -1-3 }=\frac{ -1}{ -4 }=\frac{ 1 }{ 4 }\]

OpenStudy (carson889):

Pardon me, slope is equal to rise over run. NOT run over rise.

OpenStudy (anonymous):

thanks that helps a lot